\begin{thebibliography}{9999}


% \bibitem{abrams} G. Abrams, {\it Leavitt path algebras: the first decade}, Bull. Math. Sci. 5 (2015), no. 1, 59--120.

% \bibitem{ap} G. Abrams, G. Aranda Pino, {\it The Leavitt path algebra of a graph}, J.
  %  Algebra $\mathbf{293}$ (2) (2005), 319--334.
  
\bibitem{lpabook} G. Abrams, P. Ara, M. Siles Molina, Leavitt path algebras, Lecture Notes in Mathematics {\bf 2191} Springer, 2017.

%\bibitem{abramstom} G. Abrams, M. Tomforde, {\it Isomorphism and Morita equivalence of graph algebras,} Trans. AMS {\bf 363 (7)}, 3733--3767. 


%\bibitem{aajz} A. Alahmadi, H. Alsulami, S.K. Jain and E. Zelmanov, {\it Leavitt path algebras of finite Gelfand-Kirillov dimension}, J. Algebra Appl. $\mathbf{11}$ (6) (2012), 6pp.

%\bibitem{aragoodearl} P. Ara, K. Goodearl, {\it Leavitt path algebras of separated graphs}, J. reine angew. Math. $\mathbf{669}$ (2012), 165--224.

%\bibitem{ahls} P. Ara, R. Hazrat, H. Li, A. Sims, {\it Graded Steinberg algebras and their representations}, Algebra and Number theory, $\mathbf{12-1}$ (2018), 131--172.


\bibitem{bosa} P. Ara, J. Bosa, R. Hazrat, A. Sims, {\it Reconstruction of graded groupoids from graded Steinberg algebras}, Forum Math. {\bf 29(5)} (2017), 1023--1037. 

\bibitem{pchr} G. Aranda Pino, J. Clark, A. an Huef,  I. Raeburn, {\it Kumjian--Pask
  algebras of higher-rank graphs}, Trans. Amer. Math. Soc. $\mathbf{365}$ (2013), no.
    7, 3613--3641.


%\bibitem{bergh} M. van den Bergh, {\it A note on graded K-theory}, Comm. Algebra $\mathbf{14}$ (1986), no. 8, 1561--1564.

%\bibitem{carlsenruizsims} T.M. Carlsen, E. Ruiz, A. Sims, {\it Equivalence and stable isomorphism of groupoids, and diagonal-preserving stable isomorphisms of graph $C^*$-algebras and Leavitt path algebras}, Proc. Amer. Math. Soc. $\mathbf{145}$ (2017),1581--1592.


%\bibitem{carlseneilers}  T.M. Carlsen, S. Eilers, E. Ortega, G. Restorff, {\it Flow equivalence and orbit equivalence for shifts of finite type and isomorphism of their groupoids,}  J. Math. Anal. Appl. $\mathbf{469}$ (2019), no. 2, 1088--1110. 






%\bibitem{bcw} N. Brownlowe, T.M. Carlsen, M.F. Whittaker, {\it Graph algebras and orbit equivalence}, Ergodic Theory Dynam. Systems $\mathbf{37}$ (2017), no. 2, 389--417. 





% \bibitem{bcfs} J. Brown, L.O. Clark, C. Farthing, A. Sims, {\it Simplicity of algebras
  %  associated to \'etale groupoids}, Semigroup Forum $\mathbf{88}$ (2014), no. 2, 433--452.

%\bibitem{bussexel} A. Buss, R. Exel, {\it Inverse semigroup expansions and their actions on $C^*$-algebras}, Illinois J. Math., $\mathbf{56}$ (2014), 1185--1212.

%\bibitem{clarkmichell} L.O. Clark, C. Edie-Michell, {\it Uniqueness theorems for Steinberg
   % algebras}, Algebr. Represent. Theory $\mathbf{18}$ (2015), no. 4, 907--916.

%\bibitem{clarkhazratrigby}  L.O. Clark, R. Hazrat, S.W. Rigby, {\it Strongly graded groupoids and strongly graded Steinberg algebras}, arXiv:1711.04904


\bibitem{brown} J. Brown, G.  Nagy, S. Reznikoff, A. Sims, D. Williams,  {\it Cartan subalgebras in $C^*$-algebras of Hausdorff \'etale groupoids}, 
Integral Equations Operator Theory {\bf 85} (2016), no. 1, 109--126. 


\bibitem{toke} T.M. Carlsen, E. Ruiz, A. Sims, M. Tomforde,  {\it Reconstruction of groupoids and C*-rigidity of dynamical systems}, arXiv:1711.01052.


\bibitem{carl} T.M. Carlsen, {\it $*$-isomorphism of Leavitt path algebras over $\mathbb Z$}, Adv. in Math. {\bf 324} (2018) 326--335. 

\bibitem{ccn} L. O. Clark. C. Gil Canto, A. Nasr-Isfahani, {\it The cycline subalgebra of a Kumjian--Pask algebra}, Proc. Amer. Math. Soc. $\mathbf{145}$ (2017), 1969-1980. 

\bibitem{ch}  L.O. Clark, R. Hazrat, {\it \'Etale groupoid and Steinberg algebras, a concise introduction},  arXiv:1901.01612. 


\bibitem{cep} L.O. Clark, R. Exel, E. Pardo, {\it A generalised uniqueness theorem and the graded ideal structure of Steinberg algebras}, Forum Mathematicum, $\mathbf{30}$ (2018), no. 3, 533--552.

%\bibitem{cfst} L.O. Clark, C. Farthing, A. Sims, M. Tomforde, {\it A groupoid
%    generalisation of Leavitt path algebras}, Semigroup Forum $\mathbf{89}$ (2014),
 %   501--517.

%\bibitem{cm} L.O. Clark, C. Edie-Michell, {\it Uniqueness theorems for Steinberg
 %   algebras}, Algebr. Represent. Theory $\mathbf{18}$ (2015), no. 4, 907--916.

%\bibitem{cp} L.O. Clark, Y.E.P. Pangalela, {\it Kumjian--Pask algebras of
  %  finitely-alighed higher-rank graphs}, arXiv:1512.06547v1.


%\bibitem{cs} L.O. Clark, A. Sims, {\it Equivalent groupoids have Morita equivalent
 %   Steinberg algebras}, J. Pure Appl. Algebra $\mathbf{219}$ (2015), 2062--2075.


%\bibitem{cm1984} M. Cohen, S. Montgomery, {\it Group-graded rings, smash products, and
%    group actions}, Trans. Amer. Math. Soc. $\mathbf{282}$ (1984), 237--258.



%\bibitem{crainicmoerdijk}
%M. Crainic, I. Moerdijk, {\it A homology theory for \'etale groupoids}, J. reine angew. Math. $\mathbf{521}$ (2000), 25--46. 


%\bibitem{errs} S. Eilers, G. Restorff, E. Ruiz,  A.P.W. S\o rensen, {\it Filtered K-theory for graph algebras}, 2016 MATRIX Annals, MATRIX Book Series 1,  229--249.

%\bibitem{errs2} S. Eilers, G. Restorff, E. Ruiz, A.P.W. S\o rensen, {\it The complete classification of unital graph $C^*$-algebras: Geometric and strong,}



\bibitem{gil} C. Gil Cantoa, A. Nasr-Isfahani {\it The commutative core of a Leavitt path algebra}, J. of Algebra {\bf 511} (2018) 227--248. 



%\bibitem{de} M. Dokuchaev, R. Exel, {\it Associativity of crossed products by partial actions, enveloping actions and partial representations}, Trans. Amer. Math. Soc. $\mathbf{357}$ (2005), no. 5, 1931--1952.

%\bibitem{drs} M. Dokuchaev, A. del R$\acute{\rm i}$o, J. J. Sim$\acute{\rm o}$n, {\it Globalizations of partial
%actions on non unital rings}, Proc. AMS, $\mathbf{135}$ (2007).


%\bibitem{dej} M. Dokuchaev, R. Exel, J. J. Sim$\acute{\rm o}$n, {\it Globalization of twisted
%partial actions}, Trans. AMS (2010)-----.





%\bibitem{exel2008} R. Exel, {\it Inverse semigroups and combinatorial $C^*$-algebras}, Bull. Braz. Math. Soc. (N.S.) $\mathbf{39}$ (2) (2008) 191--313.

%\bibitem{exel1994} R. Exel, {\it Circle actions on $C^*$-algebras, partial automorphisms and a generalized Pimsner-Voiculescu exact sequence}, J. Funct. Anal. $\mathbf{122}$ (1994), 361--401.


%\bibitem{exel1997} R. Exel, {\it Twisted partial actions: A classification of regular $C^*$-algebraic bundles}, Proc. London Math. Soc. $\mathbf{74}$ (3) (1997) 417--443.


%\bibitem{exel1998} R. Exel, {\it Partial Actions of Groups and Actions of Semigroups}, Proc. Am. Math. Soc. $\mathbf{126}$ (12) (1998) 3481--3494.



%\bibitem{exel} R. Exel, {\it Reconstructing a totally disconnected groupoid from its ample semigroup}, Proc. Amer. Math. Soc. $\mathbf{138}$, 2991--3001 (2010)

%\bibitem{el} R. Exel, M. Laca, {\it Cuntz-Krieger algebras for infinite matrices}, J. Reine Angew. Math. $\mathbf{512}$ (1999) 119--172.



%\bibitem{exelvieira} R. Exel, F. Vieira, {\it Actions of inverse semigroups arising from partial actions of groups}, J. Math. Anal. Appl. $\mathbf{363}$ (2010), 86--96.

%\bibitem{ferrero} M. Ferrero, {\it Partial actions of groups on algebras, a survey}, S$\tilde{\rm a}$o Paulo J. Math. Sci. $\mathbf{3}$ (2009), no. 1, 95--107.


%\bibitem{goncalvesroyer} D. Gon\c{c}alves, D. Royer, Leavitt path algebras as partial skew group rings, Comm. in Algebra $\mathbf{42}$ (2014), 3578--3592.


%\bibitem{goodearlbook} K.R. Goodearl, von Neumann regular rings, 2nd ed., Krieger Publishing Co., Malabar, FL, 1991.


%\bibitem{gr} E.L. Green, {\it Graphs with relations, coverings and group-graded
 %   algebras}, Trans. Amer. Math. Soc. $\mathbf{279}$ (1983), no. 1, 297--310.

%\bibitem{haefner} J. Haefner, {\it Graded equivalence theory with applications}, J. Algebra $\mathbf{172}$ (1995), 385--424.

%\bibitem{hazratp}  R. Hazrat and R. Preusser, {\it Applications of normal forms for weighted Leavitt path algebras: simple rings and domains},     Algebr Represent Theor (2017). 

%\bibitem{roozbehhazrat2013} R. Hazrat, {\it The graded Grothendieck group and the classication of Leavitt path algebras}, Math. Annalen $\mathbf{355}$ (2013), 273--325.



%\bibitem{grbook} R. Hazrat, Graded rings and graded Grothendieck groups,  London Math. Society Lecture Note Series $\mathbf{435}$, Cambridge University Press, 2016. 


%\bibitem{hl} R. Hazrat, H. Li, {\it Graded Steinberg algebras and partial actions}, J. Pure App. Algebra (2018), $\mathbf{222}$ (2018) 3946--3967.




%\bibitem{kp} A. Kumjian, D. Pask, {\it $C^{*}$-algebras of directed graphs and group actions}, Ergod. Th. Dynam. Systems $\mathbf{19}$ (1999), 1503--1519.
    
    
%\bibitem{lindmarcus} D. Lind and B. Marcus. An introduction to symbolic dynamics and coding, Cambridge University
%Press, 1995.


%\bibitem{matui} H. Matui, {\it Homology and topological full groups of \'etale groupoids on totally disconnected spaces}, Proc. London Math. Soc. $\mathbf{104}$ (2012), 27--56.

%\bibitem{matui15}  H. Matui, {\it Topological full groups of one-sided shifts of finite type}, J. Reine Angew. Math. $\mathbf{705}$ (2015) 35--84.

%\bibitem{matui16} H. Matui, {\it \'Etale groupoids arising from products of shifts of finite type}, Advances in Mathematics $\mathbf{303}$ (2016) 502--548. 


\bibitem{nagy} G. Nagy, S. Reznikoff, {\it Abelian core of graph algebras,} J. London Math. Soc. {\bf 85} (2012) 889--908. 

%\bibitem{maclanahan} K. McClanahan, {\it K-theory for partial crossed products by discrete groups}, J. Funct. Anal. $\mathbf{130}$ (1995) 77--117. 

%\bibitem{nastasesu-vanoystaeyen} C. N\u{a}st\u{a}sescu, F. Van Oystaeyen, Methods of Graded Rings, Lecture Notes in Math, Vol 1836. Berlin: Springer-Verlag, 2004.



%\bibitem{paterson} A.L.T. Paterson, Groupoids, inverse semigroups, and their operator algebras, BirkhaÌuser, 1999.



% \bibitem{quiggraeburn} J. C. Quigg, I.  Raeburn, {\it Characterizations of crossed products by partial actions}, J. Operator Theory $\mathbf{37}$ (1997) 311--340.


\bibitem{rr} J. Renault, {\it Cuntz-like algebras}, Operator theoretical methods. (Timioara, 1998), Theta Found (2000) 371--386. 

\bibitem{renault} J. Renault, A groupoid approach to $C^{*}$-algebras, Lecture Notes in Mathematics, {\bf 793}  Springer, Berlin, 1980.


%\bibitem{sieben} N. Sieben, {\it $C^*$-crossed products by partial actions and actions of inverse semigroups}, J. Austral. Math. Soc. Ser. A $\mathbf{63}$ (1997), 32--46.

\bibitem{st} B. Steinberg, {\it A groupoid approach to discrete inverse semigroup
    algebras}, Adv. Math. $\mathbf{223}$ (2010), 689--727.
    
\bibitem{st2} B. Steinberg algebra, {\it Diagonal-preserving isomorpisms of \'etale groupoid algebras}, J. Algebra $\mathbf{518}$ (2019), 412--439.

%\bibitem{tomforde} M. Tomforde, Uniqueness theorems and ideal structure for Leavitt path algebras, J. Algebra $\mathbf{318}$ (2007), 270--299.

%\bibitem{we} S.B.G. Webster, {\it The path space of a directed graph}, Proc. Amer. Math.
  %  Soc. $\mathbf{142}$  (2014), 213--225.



%\bibitem{xinli} X. Li, Partial transformation groupoids attached to graphs and semigroups, Int. Math.Res. Not. IMRN, to appear.


%\bibitem{wisbauer} R. Wisbauer, Foundations of module and ring theory. A handbook for study and research. Revised and translated from the 1988 German edition. Algebra, Logic and Applications, 3. Gordon and Breach Science Publishers, Philadelphia, PA, 1991. 


%\bibitem{goehle} G. Goehle, {\it Groupoid crossed products, PhD thesis}, Dartmouth College, USA, 2009.

%\bibitem{sw} A. Sims, D. P. Williams, {\it Renault's equivalence theorem for reduced groupoid $C^*$-algebras}, J. Operator Theory $\mathbf{68}$ (2012), no. 1, 223--239.

%\bibitem{aidan} A. Sims, {\it Hausdorff  \'etale groupoids and their $C^*$-algebras},  to appear in ``Operator algebras and dynamics: groupoids, crossed products and Rokhlin dimension" (F. Perera, Ed.) in Advanced Courses in Mathematics. CRM Barcelona, Birkh\"auser. 

%\bibitem{we} S.B.G. Webster, {\it The path space of a directed graph}, Proc. Amer. Math.
   % Soc. $\mathbf{142}$  (2014), 213--225.
    
    %\bibitem{tomforde} M. Tomforde, {\it Uniqueness theorems and ideal structure for Leavitt path algebras}, J. Algebra $\mathbf{318}$ (2007), 270--299.


\bibitem{we} S.B.G. Webster, {\it The path space of a directed graph}, Proc. Amer. Math.
    Soc. $\mathbf{142}$  (2014), 213--225.

    
 \end{thebibliography}

\documentclass[a4paper,reqno]{amsart}

\usepackage{graphicx}
\usepackage{mathrsfs}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amscd}
\usepackage[all,2cell]{xy}
\usepackage{color}
\usepackage[pagebackref,colorlinks]{hyperref}
\usepackage{enumerate}
\usepackage{bm}
\usepackage{bbm}
\usepackage[normalem]{ulem}
\usepackage{mathrsfs}

\usepackage{comment}
\usepackage{hyperref}
\UseAllTwocells \SilentMatrices

\numberwithin{equation}{section}

\usepackage{comment}

\theoremstyle{plain}
\newtheorem{thm}{Theorem}[section]

\newtheorem{cor}[thm]{Corollary}

\theoremstyle{definition}
\newtheorem{deff}[thm]{Definition}

\theoremstyle{remark}
\newtheorem{rmk}[thm]{\bf Remark}

\def\mG{\mathcal{G}}
\def\xra{\xrightarrow[]{}}
\def\a{\alpha}
\def\b{\beta}

\newcommand{\Isoo}{\operatorname{Iso}}
\newcommand{\spann}{\operatorname{span}}

\def\-{\text{-}}

\newcommand{\KP}{\operatorname{KP}}

\begin{document}

\title{A note on the centraliser of a subalgebra of Steinberg algebra}

\author{Roozbeh Hazrat}
\address{
Centre for Research in Mathematics and Data Sceince\\
Western Sydney University\\
Australia} \email{r.hazrat@westernsydney.edu.au}

\author{Huanhuan Li}

\address{ 1 School of Mathematical Sciences, Anhui University, Hefei, Anhui Province, China}
\address{
2 Centre for Research in Mathematics and Data Sceince\\
Western Sydney University\\
Australia} \email{h.li@westernsydney.edu.au}

\subjclass[2010]{22A22, 18B40,19D55}

\keywords{ample groupoid, Steinberg algebra, centraliser, Leavitt path algebra, diagonal of Leavitt path algebra, commutative core of Leavitt path algebra}

\date{\today}

\begin{abstract} 
For an ample Hausdorff groupoid $\mG$, and the Steinberg algebra $A_R(\mG)$ with coefficients in the commutative ring $R$ with unit,  we describe the centraliser of subalgebra $A_R(U)$ with $U$ an open closed invariant subset of unit space of $\mG$. In particular, we obtain that the algebra of  the interior of the isotropy is indeed the centraliser of the diagonal subalgebra of Steinberg algebra.  This will unify several results in the literature and  the corresponding results for Leavitt path algebras follow. 
\end{abstract}

\maketitle

\section{Introduction}

Let $E$ be a graph and $L_R(E)$  the Leavitt path algebra associated to $E$ with coefficients in the commutative ring $R$. The $R$-subalgebra of $L_R(E)$ generated by all the monomials $aa^*$, where $a$ is a path in $E$, is called the 
\emph{diagonal subalgebra} and denoted by $D_R(E)$. The \emph{commutative core} subalgebra  of $L_R(E)$ is $R$-subalgebra generated by all the monomials $aba^*$, where $a$ is a path and $b$ is a cycle with no exit and denoted by $M_R(E)$.  The commutative core subalgebra of a Leavitt path algebras were considered by Gil Cantoa and Nasr-Isfahani \cite{gil}. They prove that $M_R(E)$ is a maximal commutative subalgebra of $L_R(E)$ which coincides with  the centraliser of $D_R(E)$. They also raised an open question as to which $R$ and $E$, the Leavitt path algebra $L_R(E)$ has the property that $Z(L_R(E)) = M_R(E)$, where $Z(L_R(E))$ is the centre of $L_R(E)$. Their results are the algebraic analogue of an earlier work by Nagy and Reznikoff who introduced the $M_R(E)$ in the setting of graph $C^*$-algebras and proved that it is a masa~\cite{nagy}. A key tool in the above papers is the conditional expectation which was used in \cite{nagy} and then developed in \cite{gil} in the algebraic setting. The diagonal subalgebra $D_R(E)$ and consequently the core subalgebra $M_R(E)$ play vital roles in the structure and classification theory of graph algebras~\cite{carl}. 

In this note we consider the centraliser of a subalgebra of Steinberg algebra. In particular, we show that the core is indeed the centraliser of the diagonal subalgebra of Steinberg algebra. Despite the fact that these algebras are much more general than graph algebras, the proofs we present here are very short. Our approach does not require the use of conditional expectation and therefore it is direct and short.  Specialising to Leavitt path algebras we obtain the result of \cite{gil} and answer the open question raised there. 

\section{The centraliser of a subalgebra of Steinberg algebra}

We first recall the notions of ample groupoids and Steinberg algebras. We describe the centraliser of a subalgebra  associated to an open invariant subset of the unit space of the given groupoid. In particular, we show that the algebra of  the interior of the isotropy is indeed the centraliser of the diagonal subalgebra of Steinberg algebra; and we translate this to the settings of Leavitt path algebras and Kumjian--Pask algebras.

\subsection{Ample groupoids and Steinberg algebras}

A groupoid is a small category in which every morphism is invertible. For a groupoid $\mG$ and  $g\in\mG$,  denote $s(g):=g^{-1}g$ and $r(g):=gg^{-1}$. The pair $(g_1,g_2)$ is composable if and only if $r(g_2)=s(g_1)$. The set
$\mG^{(0)}:=s(\mG)=r(\mG)$ is called the \emph{unit space} of $\mG$.  The \emph{isotropy group} at a unit $u$ of $\mG$ is the group ${\rm Iso}(u) =\{g\in\mG \mid s(g)=r(g)=u\}$. Let ${\rm Iso}(\mG)=\bigsqcup_{u\in\mG^{(0)}}{\rm Iso}(u)$. For
$U,V\in\mG$, we define
\[
    UV=\big \{g_1 g_2 \mid g_1\in U, g_2\in V \text{ and } r(g_2)=s(g_1)\big\}.
\]

A \emph{topological groupoid} is a groupoid endowed with a topology under which the inverse map
is continuous, and such that composition is continuous with respect to the relative
topology on $\mG^{(2)} := \{(g_1,g_2) \in \mG \times \mG : s(g_1) = r(g_2)\}$ inherited from
$\mG\times \mG$. An \emph{\'etale} groupoid is a topological groupoid $\mG$ such that the
domain map $s:\mG \rightarrow \mG^{(0)}$ is a local homeomorphism. In this case, the range map $r$ is also a local
homeomorphism. Throughout this article, we assume that the \'etale groupoids are locally compact Hausdorff. 
An \emph{open bisection} of $\mG$ is an open subset $U\subseteq \mG$ such
that $s|_{U}$ and $r|_{U}$ are homeomorphisms onto an open subset of $\mG^{(0)}$. If $\mG$ is an \'etale groupoid, then there is a
base for the topology on $\mG$ consisting of open bisections with compact closure. As demonstrated in \cite{st}, if $\mG^{(0)}$ is totally disconnected and $\mG$ is \'etale, then there is a basis for the topology on $\mG$ consisting of compact open bisections. We say
that an \'etale groupoid $\mG$ is \emph{ample} if there is a basis consisting of compact open bisections for its topology. For an ample groupoid $\mG$ and a commutative ring $R$ with the discrete topology,  the $R$-value maps on $\mG$ which are locally constant and have compact support constitute an algebra called \emph{Steinberg algebra} and denoted by $A_R(\mG)$. One can show that \cite[\S 4]{ch}
\begin{align*}
A_R(\mG )& =  \Big \{\sum_{B\in F} r_B 1_B \mid  F: \text{ mutually disjoint finite collection of compact open bisections} \Big \}, \\
& \text{ addition and scalar multiplication of functions are pointwise},\\
& \text{ multiplications on the generators are }  1_B 1_D = 1_{BD}.
\end{align*}

Throughout this note we work with the Steinberg algebras associated to the ample Hausdorff groupoid $\mG$, its  unit space $\mG^{(0)}$ and its the interior of isotropy  $\Isoo(\mG)^\circ$.  

\subsection{The centraliser of a subalgebra of Steinberg algebra}

Let $\mathcal{G}$ be an ample Hausdorff groupoid. 
One can consider $A_R(\mG^{(0)})$ and $A_R(\Isoo(\mG)^\circ)$ as subalgebras of $A_R(\mG)$.  For basics of the theory of Steinberg algebras we refer the reader to \cite{ch,st}.

\begin{deff}
Let $\mG$ be an ample Hausdorff groupoid and $A_R(\mG)$ the Steinberg algebra associated to $\mG$. We call the subalgebra $A_R(\Isoo(\mG)^\circ)$ the \emph{core} algebra of $A_R(\mG)$.
\end{deff}

We need the following notation: For a subset $X$ of a ring $A$, its \emph{centraliser} (or \emph{commutant})  is defined as $C_A(X):= \{a \in A \mid xa=ax, \text{ for all }  x \in X \}$. For a subring $B\subseteq A$, we have $C_A(B)=B$ if and only if $B$ is a maximal commutative subring of $A$. 

\begin{thm}\label{maniiik}
Let $\mG$ be an ample Hausdorff groupoid and $A_R(\mG)$ the Steinberg algebra associated to $\mG$. Let $U$ be an open closed invariant subset of $ \mG^{(0)}$. Then
\begin{equation}\label{coreyy}
C_{A_R(\mG)}\big (A_R(U)\big) = A_R(\Isoo(\mG)^\circ)+A_R(\mG_{\mG^{(0)}\setminus U}).
\end{equation}
\end{thm}
\begin{proof}
Let $V \subseteq \Isoo(\mG)^\circ$ be a compact open bisection and $W \subseteq \mG^{(0)}$ a compact open set. We first observe that $WV=VW$. If $x \in WV$, then $x=wv$, where $w\in W \subseteq \mG^{(0)} $ and $v\in V$. Since $s(v)=r(v)=w$, it follows that $x=wvw=vw \in VW$. The converse follows from symmetry.  
Now for any element $x\in A_R(\Isoo(\mG)^\circ)$ and $y\in  A_R(U)$, write 
$x=\sum r_i 1_{V_i}$ and $y=\sum s_j 1_{W_j}$, where $V_i \subset \Isoo(\mG)^\circ$ and $W_j \subseteq U$ are compact open bisections. Since $V_i$ and $W_j$ commutes, we have $xy=yx$. This shows that  $A_R(\Isoo(\mG)^\circ) \subseteq C_{A_R(\mG)}\big (A_R(U)\big)  .$  Clearly that $A_R(\mG_{\mG^{(0)}\setminus U})\subseteq C_{A_R(\mG)}(A_R(U))$ as for any $f\in A_R(\mG_{\mG^{(0)}\setminus U})$ and $g\in A_R(U)$, we have $fg=0=gf$. 

Now suppose $f\in A_R(\mG)$ commutes with elements of $A_R(U)$. Write $f=\sum_{i=1}^n r_i 1_{V_i}$, where $0\not = r_i \in R$ and $V_i \subset \mG$  are disjoint compact open bisections. 
For each $V_i$, we have that $V_i^1=d^{-1}(d(V_i)\cap U)\cap V_i$ and $V_i^2=d^{-1}(d(V_i)\cap U^c)\cap V_i $ are compact with $U^c$ the completion of $U$. It follows that $1_{V_i}=1_{V_i^1}+1_{V_i^2}$. Observe that 
$1_{V_i^2}\in A_R(\mG_{\mG^{(0)}\setminus U})$. It suffices to show that for each $i$ we have $1_{V_i^1}\in A_R(\Isoo(\mG)^{o})$. That is, we need to show that $V_i^1\subseteq \Isoo(\mG)$, for all $i$'s. Suppose that there is $x \in V_k^1$ such that $s(x)\not = r(x)$. Since $\mG^{(0)}$ is Hausdorff, there is a compact open set $W \subseteq U \subseteq \mG^{(0)}$ such that $s(x)\in W$ but $r(x) \not \in W$. It follows that $x\in V_k^1W$ but $x\not \in WV_k^1$. Since $\sum_{i=1}^n r_i1_{V_i^1}\in C_{A_R(\mG)}(A_R(U))$, we have $\sum_{i=1}^n r_i1_{V_i^1}1_W=1_W \sum_{i=1}^n r_i1_{V_i^1}$. It follows that 
\begin{equation}\label{hghyty}
\Big( \sum_{i=1}^n r_i 1_{WV_i^1}\Big)(x) =\Big(\sum_{i=1}^n r_i 1_{V_i^1 W}\Big)(x).
\end{equation}
Note that $WV_1^1, \dots WV_n^1$ are disjoint and so are $V_1^1W,\dots V_n^1W$. Thus (\ref{hghyty}) reduces to $r_k=0$ which is a contradiction. So $s(x)=r(x)$. This implies that $V_i^1 \subseteq \Isoo(\mG)$ and consequently $1_{V_i^1} \in A_R(\Isoo(\mG)^\circ)$. 
\end{proof}

By Theorem \ref{maniiik} when $U=\mathcal{G}^{(0)}$, we have $C_{A_R(\mathcal{G})}(A_R(\mathcal{G}^{(0)}))=A_R(\Isoo(\mathcal{G})^o)$ which was given by \cite[Proposition 2.2]{st2}.


For the following consequence we assume that  the interior of the isotropy $\Isoo(\mG)^\circ$, is abelian (that is, it is a bundle of abelian groups). A large class of topological groupoids have abelian isotropy, such as Deaconu-Renault groupoids~\cite{rr} (and in particular graph groupoids \S\ref{hfyhf}). 

\begin{cor}\label{colkfhgy}
Let $\mG$ be an ample Hausdorff groupoid and $A_R(\mG)$ the Steinberg algebra associated to $\mG$. Suppose $\Isoo(\mG)^\circ$ is abelian. Then we have the followings. 

\begin{enumerate}[\upshape(i)]

\item The algebra $A_R(\Isoo(\mG)^\circ)$ is a maximal commutative subalgebra of $A_R(\mG)$.

\medskip 

\item The Steinberg algebra $A_R(\mG)$ is commutative if and only if $Z(A_R(\mG))=C_{A_R(\mG)}\big (A_R(\mG^{(0)}\big)$.

\end{enumerate}
\end{cor}
\begin{proof}

(1) Since $\Isoo(\mG)^\circ$ is abelian, the subalgebra $A_R(\Isoo(\mG)^\circ)$ is commutative. Now the results follows from Theorem~\ref{maniiik}.

(2) Suppose  $Z(A_R(\mG))=C_{A_R(\mG)}\big (A_R(\mG^{(0)}\big)$. Let $V$ be a compact open bisection of $\mG$. Suppose that there is $x \in V$ such that $s(x)\not = r(x)$. Since $\mG^{(0)}$ is Hausdorff, there is a compact open set $U \subseteq \mG^{(0)}$ such that $s(x)\in U$ but $r(x) \not \in U$. It follows that $x\in VU$ but $x\not \in UV$. But by assumption, $1_U \in A_R(\mG^{(0)}) \subseteq Z(A_R(\mG))$ and thus $1_{UV}=1_U 1_V = 1_V 1_U= 1_{VU}$, i.e., $UV=VU$ which is a contradiction. Thus $V \subseteq \Isoo(\mG)$.  So $A_R(\mG)=A_R(\Isoo (\mG)^{o})$, and therefore $A_R(\mG)$ is commutative.  The converse is clear. 
\end{proof}

If $\mG$ is effective, then $\mG^{(0)}=\Isoo(\mG)^\circ$, therefore by Corollary~\ref{colkfhgy}(i), $A_R(\mG^{(0)})$ is a maximal commutative subalgebra of $A_R(\mG)$. Since topologically principal is a stronger condition than effectiveness, we obtain~\cite[Lemma 2.1]{bosa} from Corollary~\ref{colkfhgy}

\subsection{Commutative cores of Leavitt path algebras and Kumjian--Pask algebras}\label{hfyhf}
Let $E=(E^0,E^1,s,r)$ be a directed graph. We refer the reader to \cite{lpabook} for the basics on the theory of Leavitt path algebras. We denote by $E^{\infty}$ the set of
infinite paths in $E$. Set
$
X := E^{\infty}\cup  \{\mu\in E^{*}   \mid   r(\mu) \text{ is not a regular vertex}\}.
$
Let
\[
\mG_{E} := \big \{(\a x,|\a|-|\b|, \b x) \mid   \a, \b\in E^{*}, x\in X, r(\a)=r(\b)=s(x)\big \}.
\]
We view each $(x, k, y) \in \mG_{E}$ as a morphism with range $x$ and source $y$. The
formulas $(x,k,y)(y,l,z)= (x,k + l,z)$ and $(x,k,y)^{-1}= (y,-k,x)$ define composition
and inverse maps on $\mG_{E}$ making it a groupoid with $\mG_{E}^{(0)}=\{(x, 0, x) \mid
x\in X\}$ which we identify with the set $X$.
Next, we describe a topology on $\mG_{E}$. For $\mu\in E^{*}$ define
\[
Z(\mu)= \{\mu x \mid x \in X, r(\mu)=s(x)\}\subseteq X.
\]
For $\mu\in E^{*}$ and a finite $F\subseteq s^{-1}(r(\mu))$, define
\[
Z(\mu\setminus F) = Z(\mu) \setminus \bigcup_{\a\in F} Z(\mu \a).
\]
The sets $Z(\mu\setminus F)$ constitute a basis of compact open sets for a locally
compact Hausdorff topology on $X=\mG_{E}^{(0)}$ (see \cite[Theorem 2.1]{we}).
For $\mu,\nu\in E^{*}$ with $r(\mu)=r(\nu)$, and for a finite $F\subseteq E^{*}$ such
that $r(\mu)=s(\a)$ for $\a\in F$, we define
\[
Z(\mu, \nu)=\{(\mu x, |\mu|-|\nu|, \nu x) \mid  x\in X, r(\mu)=s(x)\},
\]
and then
\[
Z((\mu, \nu)\setminus F) = Z(\mu, \nu) \setminus \bigcup_{\a\in F}Z(\mu\a, \nu\a).
\]
The sets $Z((\mu, \nu)\setminus F)$ constitute a basis of compact open bisections for a
topology under which $\mG_{E}$ is an ample groupoid. We have an isomorphism 
the map
\begin{align*}
\pi_{E} : L_{R}(E) &\longrightarrow A_{R}(\mG_{E})\\
v &\longmapsto 1_{Z(v)},\\
e &\longmapsto 1_{Z(e, r(e))},\\
e^* &\longmapsto 1_{Z(r(e), e)},
\end{align*} 
where $v\in E^{0}$ and $e\in E^{1}$. Recall that the diagonal subalgebra $D_R(E)$ is generated by monomials $aa^*$, where $a$ is a path in $E$, whereas the commutative core subalgebra  $M_R(E)$ is generated the monomials $aba^*$, where $a$ is a path and $b$ is a cycle with no exit. 
Observe that the isomorphism $\pi_E$ restricts to isomorphisms 
\begin{align}\label{assteinberg}
D_R(E) &\longrightarrow A_R(\mG^{(0)}) \\
M_R(E) &\longrightarrow  A_R(\Isoo(\mG)^{\circ}). \notag
\end{align}

We can recover the results of \cite{gil} and answer the open question raised there \cite[p. 245]{gil}. 

\begin{cor}
Let $E$ be an arbitrary graph and $L_R(E)$ the Leavitt path algebra associated to $E$. Then 

\begin{enumerate}[\upshape(i)]

\item The centraliser of the diagonal algebra $D_R(E)$ is the core algebra $M_R(E)$. 

\item The core algebra $M_R(E)$ is the maximal commutative subalgebra of $L_R(E)$. 

\item If $Z(L_R(E))= M_R(E)$ then $L_R(E)$ is either $R$ or $R[x,x^{-1}]$, i.e, the graph $E$ is either a single vertex or a vertex with a loop. 

\end{enumerate}
\end{cor}
\begin{proof}

(i) and (ii) immediately follow follow from Theorem~\ref{maniiik} and Corollary~\ref{colkfhgy} by considering the graph groupoid $\mG_E$ (see \ref{assteinberg}). 

(iii) By Corollary~\ref{colkfhgy}, $L_R(E)$ is commutative. Thus the graph $E$ is either a single vertex or a vertex with a loop. 
\end{proof}

In a similar manner we can determine the centraliser of the diagonal algebra of Kumjian-Pask algebras as they can also be described as a Deaconu-Renault groupoid algebras. 
Let $\Lambda$ be a row-finite $k$-graph without sources and $\KP_K(\Lambda)$ the
Kumjian--Pask algebra of $\Lambda$. We refer the reader to \cite{pchr} for the basics on the theory Kumjian-Pask algebras.  Following \cite{pchr}, $\Lambda^\infty$ denotes the set of all degree-preserving functor $x : \Omega_{k}\xra \Lambda$. Here  $\Omega_k$ is the $k$-graph defined as a set by $\Omega_k = \{(m,n)
\in \mathbb{N}^k \times \mathbb{N}^k : m \le n\}$ with $d(m,n) = n-m$, $\Omega_k^0 =
\mathbb{N}^k$, $r(m,n) = m$, $s(m,n) = n$ and $(m,n)(n,p) = (m,p)$.

The \emph{diagonal algebra} $D_R(\Lambda)$ is generated by $\lambda \lambda^*$ where $\lambda \in \Lambda$. We define the \emph{core} subalgebra 
\[M_R(\Lambda):=\spann_R \, \big \langle \lambda \mu^* \mid  \lambda x = \mu x, \text{ for any } x\in \Lambda^{\infty} \big\rangle.\]

By appealing to  Theorem~\ref{maniiik} and Corollary~\ref{colkfhgy} , similar to the case of Leavitt path algebras, we can recover the results of \cite[Theorem 4.6]{ccn}. 
\begin{cor}\label{kpk}
Let $\Lambda$ be a row-finite $k$-graph with no sources and  $\KP_K(\Lambda)$ the
Kumjian--Pask algebra associated to $\Lambda$. Then 

\begin{enumerate}[\upshape(i)]

\item The centraliser of the diagonal algebra $D_R(\Lambda)$ is the core algebra $M_R(\Lambda)$. 

\item The core algebra $M_R(\Lambda)$ is the maximal commutative subalgebra of $\KP_R(\Lambda)$. 

\item If $Z(\KP_R(\Lambda))= M_R(\Lambda)$ then $\KP_R(\Lambda)$ is $R[x_1^{\pm 1},  \dots, x_k^{\pm 1}]$. 

\end{enumerate}
\end{cor}

\begin{rmk}
In the setting of groupoid $C^*$-algebras, the general version of Theorem~\ref{maniiik} yet to be established. In this setting, with an extra assumption that $\Isoo(\mG)^\circ$ is abelian, the $C^*$-version of Theorem~\ref{maniiik} was established in \cite{toke} (see \cite[Corollary 5.3, 5.4]{toke}), and also \cite[Theorem~4.3]{brown}). The latter Theorem was used to give the $C^*$-version of Corollary~\ref{kpk} (i) and (ii) (see \cite[Corollary 4.6]{brown}). 
We finish the note by remarking that Exel, Clark and Pardo~\cite{cep} have established the algebraic version of uniqueness theorem of \cite[Theorem 3.1]{brown}. Namely, let $\mG$ be a second-countable, ample, Hausdorff groupoid, $R$ a unital commutative ring, and let $\pi: A_R(\mG) \rightarrow A$ be a homomorphism of rings.  Then $\pi$ is injective if and only if the restriction of $\pi$ on the core subalgebra $A_R(\Isoo(\mG)^{\circ})$ is injective. 

\end{rmk}

\section{Acknowledgements} The authors would like to acknowledge Australian Research Council grant DP160101481. They would like to thank Aidan Sims for fruitful discussions and for pointing them to the papers \cite{brown,toke}.

\end{document}